How to find magnetic moment of semicircle when length of wire and pole strength are given?


question is:- 62.8cm long magnetized wire is bent into the form of a semicircle.The pole strength is 2Am.Moment of resultant magnet is?


a)2*31.4


b)2*.314


c)2* .1


d)2*.4


e)2*.628

Grade:12

2 Answers

Shyama P
12 years ago

Effective length =2R=2l/pi

Mag. moment= p*2R= 2*62.8*2*10-2/3.14

Is the ans: 2*.4

Sanandanan M
12 years ago

yeah, thanks shyama